I have a question which goes like this:
An = 2-A*n-1, where A5 = 3. I need to find the tenth value. The answer is -1 but I don't know exactly how to do it.
Can you please help me out?
Thank you

Answer by rothauserc(4718)

we are given
An = 2-A*n-1, where A5 = 3, what is A10?
we have
3 = 2-A*n-1
there are no parenthesis so we follow the order of operations (do multiplication and division from left to right in the order they appear, then addition and subtraction from left to right in the order they appear)
A5=3 = 2 - 5A - 1
3+5A = 1
5A = -2
A = -2/5
now
A10 = 2-((-2/5)*10)-1
A10 = 2-(-4)-1
A10 = 2+4-1
A10 = 5